OpenStudy (anonymous):

Express eqyation, x as a co of 7 func u and t

myininaya (myininaya):

Notice wherever you have a variable, it is x^2 So what do you think the outside function and inside function could be?

myininaya (myininaya):

Example: Say we have F(x)=1+x^3+x^6 And we wanted to identify functions f(x) and g(x) such that F(x)=f(g(x)) . Well F(x)=1+x^3+(x^3)^2 Wherever you see a variable, it is x^3 so inside function is x^3, in other words let g(x)=x^3 Now replace x^3 with x and you have your outside function f(x)=1+x+x^2 I hope that is a pretty clear example